# Boosting In-Silicon Directed Evolution with Fine-Tuned Protein Language Model and Tree Search

### v1

@@ Abstract
Protein evolution through amino acid sequence mutations is a cornerstone of life sciences. While current in-silicon directed evolution algorithms focus on designing search strategies, they overlook how to utilize the transformative protein language models, which encode rich evolutionary patterns, to guide search. To bridge this gap, we propose AlphaDE, a novel framework to evolve protein sequences by harnessing the innovative paradigms of large language models. First, AlphaDE fine-tunes pretrained protein language models using masked language modeling on homologous protein sequences to activate the evolutionary plausibility for the interested protein class. Second, AlphaDE introduces test-time inference based on Monte Carlo tree search, which effectively evolves proteins with evolutionary guidance from the fine-tuned protein language model. Extensive benchmark experiments show that AlphaDE remarkably outperforms previous state-of-the-art methods even with few-shot fine-tuning. An interesting case study further shows that AlphaDE supports condensing the protein sequence space through computational evolution.
@@ 1 Introduction
Proteins are essential components of living systems, exhibiting a great diversity of functions among biological macromolecules (holm_mapping_1996). They play critical roles in a wide range of biochemical processes, including enzyme catalysis, cellular metabolism, immune responses, and signal transduction (jiang_general_2024). Protein engineering through directed evolution enables optimization of protein functions by seeking potential protein variants with improved properties such as the expression level and catalytic activity (yang_machine-learning-guided_2019). Traditional in vitro or in vivo experimental directed evolution approaches, such as deep mutational scanning (fowler_deep_2014) and orthogonal DNA replication system (ravikumar_orthogonal_2014), directly measure the functional effects of protein mutations but are limited to exploring only a fraction of the possible protein space, and are usually expensive and laborious. To reduce the burden of the expensive wet experiments, recent advances in machine learning-guided approaches focus on building a surrogate sequence-function landscape (yang_machine-learning-guided_2019), and a lot of in-silicon directed evolution algorithms (brookes_design_2018; sinai_adalead_2020; ren_proximal_2022) have been proposed to strategically explore the protein fitness landscape to identify optimal sequence mutations via an iterative search process.
@@
On the other hand, protein language models (meier_language_2021; lin_evolutionary-scale_2023), which encapsulate millions of years of evolutionary information through unsupervised pretraining in massive protein databases, are rapidly changing the domain of protein design (madani_large_2023; jiang_rapid_2024; hayes_simulating_2025). As protein language models implicitly learn complex evolutionary and structural dependencies from natural protein sequences, researchers have been increasingly employing protein language models for protein engineering tasks such as zero-shot inference of the functional effects of sequence substitutions to find high-fitness variations (hie_efficient_2024; shanker_unsupervised_2024). However, despite protein language models encoding rich evolutionary clues, few in-silicon directed evolution approaches study how to leverage protein language models into directed evolution for effective exploration in the immense protein sequence space.
@@
To fill this gap, in this work, we propose a novel framework named AlphaDE to directly evolve protein sequences following the technical paradigm of natural large language models. Specifically, AlphaDE consists of a fine-tuning step and a Monte Carlo tree search (MCTS) inference step built on the protein language model. In the fine-tuning step, AlphaDE fine-tunes the pretrained protein language model with homologous protein sequences to activate its evolutionary plausibility for the interested protein class. In the MCTS inference step, similar to test-time MCTS in large language models to boost reasoning (zhou_language_2024; guan_rstar-math_2025), AlphaDE conducts an iterative tree search to efficiently optimize protein function via residue mutations guided by the fine-tuned protein language model. Through the two synergic steps, AlphaDE harvests superior directed evolution ability for proteins. To evaluate AlphaDE, we conduct computational benchmark experiments on eight distinct tasks. Impressively, with the same query budget of the simulated fitness oracle, AlphaDE substantially outperforms various in-silicon directed evolution methods. Further few-shot fine-tuning experiments reveal that AlphaDE’s evolution ability can be activated by fine-tuning with dozens of homologous sequences. Lastly, to show the broader applicability of AlphaDE, we conduct a proof-of-concept task to computationally condense the sequence space of the functional avGFP protein.
@@ 2.2 Protein Directed Evolution
Directed evolution is a classical paradigm for protein sequence design, where a plenty of algorithms are developed to accelerate the in-silicon directed evolution process. AdaLead (sinai_adalead_2020) is an advanced implementation of model-guided directed evolution with iteratively recombined and mutated operations for seed sequences. CMA-ES (hansen_completely_2001) is a second-order evolutionary search algorithm that estimates the covariance matrix to adaptively adjust the search strategy of the upcoming generations. Bayesian optimization (BO) (snoek_practical_2012) is a classical paradigm for the sequential design problem (mockus_bayesian_1989), which estimates the uncertainty and constructs the acquisition function for exploration. DbAS (brookes_design_2018) establishes a probabilistic framework that trains a variational autoencoder (VAE) (kingma_auto-encoding_2022) to model the distribution of high-fitness sequences and adaptively samples sequences from this trained VAE to explore the fitness landscape. Following DbAS, CbAS (brookes_conditioning_2019) estimates the probability distribution conditioned on the desired properties with model-based adaptive sampling and additionally considers a regularization to stabilize the model-guided search process. DyNA-PPO (angermueller_model-based_2020) formulates protein sequence design as a sequential decision-making problem and uses proximal policy optimization (PPO) (schulman_proximal_2017) to perform sequence generation. PEX (ren_proximal_2022) aims to search for effective candidates of low-order mutants near the wild-type, and formulates this process as a proximal optimization problem to solve. EvoPlay (wang_self-play_2023) uses the self-play reinforcement learning inspired by AlphaZero (silver_general_2018) to optimize protein sequences. More recently, TreeNeuralTS and TreeNeuralUCB (qiu_tree_2024) combine tree search with bandit machine learning for directed evolution tasks, which expands a tree starting from the initial sequence with the guidance of a bandit machine learning model. Specifically, TreeNeuralTS adopts Thompson sampling (thompson_likelihood_1933) while TreeNeuralUCB utilizes the upper confidence bound (auer_finite-time_2002) to explore the sequence space. Different from previous works, in this paper, we explore using the natural large language model technique paradigm to boost in-silicon directed evolution with protein language models.
@@ 3 AlphaDE
In this section, we explain the proposed AlphaDE for protein directed evolution in detail. First, we give the problem definition of the directed evolution in Section 3.1. Second, in Section 3.2, we introduce how to fine-tune the pretrained protein language model to give prior guidance on the next mutation residues. Third, in Section 3.3, we show how to perform tree search to directly evolve protein sequences following the prior mutation guidance from the fine-tuned protein language model. The framework is shown in Figure 1.
@@ 3.3 Monte Carlo Tree Search with Fine-Tuned Protein Language Model
We utilize Monte Carlo tree search (MCTS) to solve the MDP of directed evolution. Previous MCTS work (rosin_multi-armed_2011; silver_mastering_2016) trains neural networks based on expert data for experienced guidance of action selection. Similarly, to enable efficient guidance of protein mutation, AlphaDE employs the fine-tuned protein language model to predict the next mutation action (the residue position and residue type) given the current protein sequence. At each step, the fine-tuned protein language model receives the protein sequence and outputs the probability logits of each residue type on each residue position as guidance for MCTS.
When mutating protein sequences with the fine-tuned protein language model, although we could mutate protein sequences in a greedy manner by taking the next mutation with the maximum probability, it is prone to being stuck in a local optimum due to the unpredictable complexity of the protein fitness landscape (romero_exploring_2009). Moreover, the predicted mutation with the maximum probability does not mean that it is optimal because the protein language model is not optimized for a specific protein fitness. To directly maximize a given protein fitness, we integrate MCTS into AlphaDE to solve the MDP of directed evolution with the help of a fine-tuned protein language model for evolution guidance of mutation actions. Next, we introduce the MCTS part of AlphaDE in the background of protein directed evolution.
MCTS (browne_survey_2012) adopts a tree structure to perform simulation iterations and estimate the state value of actions. Meanwhile, it uses the previously estimated action values to guide the search process towards higher rewards. The MCTS part in AlphaDE consists of the following four steps per iteration:
@@
  • Rollout. The value of the reached leaf node $a_{\tau+l}$ is evaluated by a rollout. From the leaf node, MCTS recursively mutates the sequence until termination, then the final sequence $s_{T}$ is evaluated by a fitness oracle for reward $r$. During the rollout, each mutation is selected greedily. To speed up the computationally expensive rollout process, AlphaDE trains a value network to predict the state value of the reached leaf nodes as AlphaZero (silver_general_2018). Specifically, if the reached leaf node $s_{\tau+l}$ is a termination node, it is evaluated by the fitness oracle. Otherwise, it is predicted by the value network as shown in Figure 1.
@@ 4 Experiments
In this section, we conduct computational experiments to validate the effectiveness of AlphaDE. First, in Section 4.1, we benchmark AlphaDE with various directed evolution algorithms on extensive tasks. Second, we study AlphaDE under different homology sequence availability settings, such as few-shot fine-tuning in Section 4.2 and zero-shot scaling with model sizes in Section 4.3. Finally, we show that AlphaDE could be utilized to computationally condense the protein sequence space of the fluorescent protein in Section 4.4.
@@ 4.1 Benchmark experiments
We compare AlphaDE with various baselines including TreeNeuralTS, TreeNeuralUCB, PEX, EvoPlay, AdaLead, DyNA-PPO, DbAS, CbAS, BO, and CMA-ES. We follow the officially released codebase and configurations of TreeNeuralTS, TreeNeuralUCB, and EvoPlay for running. We take the results of PEX, AdaLead, CbAS, DbAS, and DyNA-PPO from Ren et al. (ren_proximal_2022), which utilizes the implementations from the FLEXS environment (sinai_adalead_2020) for evaluation, and the results are averaged over 80 trials. We run BO and CMA-ES with the implementations in FLEXS. We average the results of AlphaDE, TreeNeuralTS, TreeNeuralUCB, EvoPlay, BO, and CMA-ES over 5 independent trials. For all the baselines, the generation process is divided into 10 rounds, with 100 queries to the oracle in each round. For AlphaDE, we also set the same oracle budget of 1000 queries. Here we use the pretrained ESM2-series models (lin_evolutionary-scale_2023) for fine-tuning as it is one of the most powerful and widely used protein language models with various model sizes for study convenience. In the standard version of AlphaDE, we use ESM2-35M for fine-tuning as its efficiency in model size and effectiveness in evolution plausibility. Details of the homologous sequence datasets for fine-tuning are in Appendix B, while hyperparameters of AlphaDE are given in Appendix C. For each task, the starting protein sequence for optimization is the sequence with the lowest protein fitness value. The benchmark results for these eight tasks are summarized in Table 1. The results in Table 1 show that AlphaDE significantly outperforms various baselines in most tasks. For example, in the task of TEM, the second best method TreeNeuralTS obtains a fitness value of 0.27 while AlphaDE achieves 1.22 with an improvement of 351.85%. Similar situations also happen in the tasks of AAV, E4B, and AMIE, which exhibit AlphaDE’s great evolution ability. At the same time, we also examine the diversity of evolved sequences by AlphaDE in Appendix I. AlphaDE with different ESM2 model sizes is tested in Appendix D. The ablation study of AlphaDE is given in Appendix F. Benchmark results using another oracle based on ESM-1b further validate AlphaDE as shown in Appendix G. Moreover, the study of hyperparameter sensitivity is available in Appendix H, while AlphaDE’s computation efficacy is provided in Appendix L.
Table 1: Benchmarking in-silicon directed evolution methods. We present the maximum fitness in 1000 black-box oracle queries. Results of AlphaDE, TreeNeuralTS, TreeNeuralUCB, EvoPlay, BO, and CMA-ES are averaged over 5 independent trials while others are from Ren et al. (ren_proximal_2022).
@@
At the same time, we also test AlphaDE with other protein language models such as ProtBert (elnaggar_prottrans_2022) (model size 420M) and ESM-1b (rives_biological_2021) (model size 650M), which are trained with the standard BERT (devlin_bert_2019) architecture using the UniRef100 (suzek_uniref_2007) database and with the optimized RoBERTa (liu_roberta_2019) architecture using the UniRef50 (suzek_uniref_2007) database, respectively. The results of AlphaDE with ProtBert and ESM-1b are shown in Table 2. We see that AlphaDE with ProtBert and AlphaDE with ESM-1b perform differently in some tasks, suggesting their evolution ability varies in different proteins. Generally, AlphaDE with the two different protein language models also achieves superior performance. Meanwhile, the fine-tuning step is the key to achieving a performance leap, as directly using the pretrained versions of protein language models leads to much fitness loss.

### v2

@@ Abstract
Protein evolution through amino acid sequence mutations is a cornerstone of life sciences. While current in-silicon directed evolution algorithms largely focus on designing heuristic search strategies, they overlook how to integrate the transformative protein language models, which encode rich evolutionary patterns, with reinforcement learning to learn to directly evolve proteins. To bridge this gap, we propose AlphaDE, a novel framework to optimize protein sequences by harnessing the innovative paradigms of large language models such as fine-tuning and test-time inference. First, AlphaDE fine-tunes pretrained protein language models using masked language modeling on homologous protein sequences to activate the evolutionary plausibility for the interested protein class. Second, AlphaDE introduces test-time inference based on Monte Carlo tree search, which effectively evolves proteins with evolutionary guidance from the fine-tuned protein language model. Extensive benchmark experiments show that AlphaDE remarkably outperforms previous state-of-the-art methods even with few-shot fine-tuning. A further case study demonstrates that AlphaDE supports condensing the protein sequence space of avGFP through computational evolution.
@@ 1 Introduction
Proteins are essential components of living systems, exhibiting a great diversity of functions among biological macromolecules (holm_mapping_1996). They play critical roles in a wide range of biochemical processes, including enzyme catalysis, cellular metabolism, immune responses, and signal transduction (jiang_general_2024). Protein engineering through directed evolution enables optimization of protein functions by seeking potential protein variants with improved properties such as the expression level and catalytic activity (yang_machine-learning-guided_2019). Traditional in vitro or in vivo experimental directed evolution approaches, such as deep mutational scanning (fowler_deep_2014) and orthogonal DNA replication system (ravikumar_orthogonal_2014), directly measure the functional effects of protein mutations but are limited to exploring only a fraction of the possible protein space, and are usually expensive and laborious. To reduce the burden of the expensive wet experiments, recent advances in machine learning-guided approaches focus on building a surrogate sequence-function landscape (yang_machine-learning-guided_2019), and a lot of in-silicon directed evolution algorithms (brookes_design_2018; sinai_adalead_2020; ren_proximal_2022; wang_self-play_2023) have been proposed to strategically explore the protein fitness landscape to identify optimal sequence mutations via an iterative search process.
@@
On the other hand, protein language models (meier_language_2021; lin_evolutionary-scale_2023), which encapsulate millions of years of evolutionary information through unsupervised pretraining in massive protein databases, are rapidly changing the domain of protein design (madani_large_2023; jiang_rapid_2024; hayes_simulating_2025). As protein language models implicitly learn complex evolutionary and structural dependencies from natural protein sequences, researchers have been increasingly employing protein language models for protein engineering tasks such as zero-shot inference of the functional effects of sequence substitutions to find high-fitness variations (hie_efficient_2024; shanker_unsupervised_2024). More recently, some pioneering works have begun to introduce pretrained protein language models in the process of directed evolution (jiang_rapid_2024; yang_active_2025; tran_protein_2025). However, although employing protein language models to recommend mutations, these works use simple heuristic search methods such as greedy selection (jiang_rapid_2024) or beam search (tran_protein_2025) to search in directed evolution, leaving a large room to explore how to integrate protein language models into advanced optimization techniques such as reinforcement learning.
@@
To fill this gap, in this work, we propose a novel framework named AlphaDE to directly evolve protein sequences following the technical paradigm of natural large language models (guo_deepseek-r1_2025). Specifically, AlphaDE consists of a fine-tuning step and a Monte Carlo tree search (MCTS) inference step built on the protein language model. In the fine-tuning step, AlphaDE fine-tunes the pretrained protein language model with homologous protein sequences to activate its evolutionary plausibility for the interested protein class. In the MCTS-assisted inference step, similar to test-time MCTS in large language models to boost reasoning (zhou_language_2024; guan_rstar-math_2025), AlphaDE conducts an iterative tree search to efficiently optimize protein function via residue mutations guided by the fine-tuned protein language model. Through the two synergic steps, AlphaDE harvests superior directed evolution ability for proteins. To evaluate AlphaDE, we conduct computational benchmark experiments on eight distinct tasks. Impressively, AlphaDE substantially outperforms various in-silicon directed evolution methods. Further few-shot fine-tuning experiments reveal that AlphaDE’s evolution ability can be activated by fine-tuning with dozens of homologous sequences. Lastly, to show the broader applicability of AlphaDE, we conduct a proof-of-concept task to computationally condense the sequence space of the functional avGFP protein.
@@ 2.1 Protein Language Models
At the same time, some works (widatalla_aligning_2024; yang_steering_2025) utilize labeled sequence-fitness pairs to steer protein language models for de novo protein sequence design to generate novel and high-quality sequences, showing the superior alignment ability of protein language models.
@@ 2.2 Protein Directed Evolution
Directed evolution is a classical paradigm for protein sequence design, where a plenty of algorithms are developed to accelerate the in-silicon directed evolution process. AdaLead (sinai_adalead_2020) is an advanced implementation of model-guided directed evolution with iteratively recombined and mutated operations for seed sequences. CMA-ES (hansen_completely_2001) is a second-order evolutionary search algorithm that estimates the covariance matrix to adaptively adjust the search strategy of the upcoming generations. Bayesian optimization (BO) (snoek_practical_2012) is a classical paradigm for the sequential design problem (mockus_bayesian_1989), which estimates the uncertainty and constructs the acquisition function for exploration. DbAS (brookes_design_2018) establishes a probabilistic framework that trains a variational autoencoder (VAE) (kingma_auto-encoding_2022) to model the distribution of high-fitness sequences and adaptively samples sequences from this trained VAE to explore the fitness landscape. Following DbAS, CbAS (brookes_conditioning_2019) estimates the probability distribution conditioned on the desired properties with model-based adaptive sampling and additionally considers a regularization to stabilize the model-guided search process. DyNA-PPO (angermueller_model-based_2020) formulates protein sequence design as a sequential decision-making problem and uses proximal policy optimization (PPO) (schulman_proximal_2017) to perform sequence generation. PEX (ren_proximal_2022) aims to search for effective candidates of low-order mutants near the wild-type, and formulates this process as a proximal optimization problem to solve. EvoPlay (wang_self-play_2023) uses the self-play reinforcement learning inspired by AlphaZero (silver_general_2018) to optimize protein sequences. Recently, TreeNeuralTS and TreeNeuralUCB (qiu_tree_2024) combine tree search with bandit machine learning for directed evolution, which expands a tree starting from the initial sequence with the guidance of a bandit machine learning model. Specifically, TreeNeuralTS adopts Thompson sampling (thompson_likelihood_1933) while TreeNeuralUCB utilizes the upper confidence bound (auer_finite-time_2002) to explore the sequence space.
More recently, with the emergence and development of protein language models, some pioneering works explore to integrate the pretrained protein language models in the process of directed evolution (jiang_rapid_2024; yang_active_2025; tran_protein_2025). For example, jiang_rapid_2024 and yang_active_2025 present the multi-round active learning to rapidly improve protein activity, which alternate between collecting sequence fitness using a wet-lab assay and training a protein language model to prioritize new sequences to screen in the next round. Meanwhile, LatentDE (tran_latentde_2025) uses gradient-based method to search in the latent space represented by VAE with protein embeddings extracted from a frozen, general-purpose pretrained protein language model (i.e., ESM-2). Moreover, MLDE (tran_protein_2025) introduces an optimization pipeline that utilizes protein language models to pinpoint the mutation hotspots and then suggest replacements by heuristically calculating a k-mer’s relevancy and entropy in a sequence. However, despite leveraging protein language models to recommend mutations, these works use simple search methods such as greedy selection (jiang_rapid_2024) or beam search (tran_protein_2025), remaining how to integrate protein language models into advanced optimization techniques such as reinforcement learning unexplored in the directed evolution field. To bridge this gap, here we follow a principle of effective searching (i.e., AlphaZero-like MCTS) with strong prior guidance (i.e., fine-tuned protein language model), inspired by the natural large language model technique paradigm such as fine-tuning (shao_deepseekmath_2024) and test-time inference (guan_rstar-math_2025), to boost in-silicon directed evolution with a novel technique paradigm.
@@ 3 AlphaDE
In this section, we show AlphaDE for protein directed evolution in detail. First, we give the problem definition of the directed evolution in Section 3.1. Second, in Section 3.2, we introduce how to fine-tune the pretrained protein language model to give prior guidance on the next mutation residues. Third, in Section 3.3, we show how to perform tree search to directly evolve protein sequences following the prior mutation guidance from the fine-tuned protein language model. The whole framework of AlphaDE is shown in Figure 1.
@@ 3.3 Monte Carlo Tree Search with Fine-Tuned Protein Language Model
We use Monte Carlo tree search (MCTS) to solve the MDP of directed evolution. To effectively select actions, some MCTS works (rosin_multi-armed_2011; silver_mastering_2016) train neural networks based on expert data for experienced guidance. Similarly, to enable efficient guidance of protein mutation, AlphaDE employs a fine-tuned protein language model to predict the next mutation action (the residue position and residue type) given the current protein sequence. At each step, the fine-tuned protein language model receives the protein sequence and outputs the probability logits of each residue type on each residue position as guidance for MCTS.
When mutating protein sequences with the fine-tuned protein language model, although we could mutate a protein sequence in a greedy manner by taking the next mutation with the maximum probability, it is prone to being stuck in a local optimum due to the unpredictable complexity of the protein fitness landscape (romero_exploring_2009). Moreover, the predicted mutation with the maximum probability does not mean that it is optimal because the protein language model is not optimized for a specific protein fitness. To directly maximize a given protein fitness, we integrate MCTS into AlphaDE to solve the MDP of directed evolution with the help of a fine-tuned protein language model for mutation guidance. Next, we introduce the MCTS part of AlphaDE in the background of protein directed evolution.
MCTS (browne_survey_2012) adopts a tree structure to perform simulation iterations and estimate the state value of actions. Meanwhile, it uses the previously estimated action values to guide the search process towards higher rewards. As shown in Figure 1, the MCTS part in AlphaDE consists of the following four steps per iteration:
@@
  • Rollout. The value of the reached leaf node $a_{\tau+l}$ is evaluated by a rollout. From the leaf node, MCTS recursively mutates the sequence until termination, then the final sequence $s_{T}$ is evaluated by a fitness oracle for reward $r$. During the rollout, each mutation is selected greedily. To speed up the computationally expensive rollout process, AlphaDE trains a value network to predict the state value of the reached leaf nodes as AlphaZero (silver_general_2018). Specifically, if the reached leaf node $s_{\tau+l}$ is a termination node, it is evaluated by the fitness oracle. Otherwise, it is predicted by the value network.
@@ 4 Experiments
In this section, we conduct computational experiments to validate the effectiveness of AlphaDE. First, in Section 4.1, we benchmark AlphaDE with various directed evolution algorithms on extensive tasks. Second, we study AlphaDE under different settings of homology sequence availability, such as few-shot fine-tuning in Section 4.2 and zero-shot scaling with model sizes in Section 4.3. Finally, we utilize AlphaDE to computationally condense the protein sequence space of the fluorescent protein in Section 4.4. The ablation study of AlphaDE is provided in Appendix F.
@@ 4.1 Benchmark experiments
We compare AlphaDE with various baselines including LatentDE, MLDE, TreeNeuralTS, TreeNeuralUCB, PEX, EvoPlay, AdaLead, DyNA-PPO, DbAS, CbAS, BO, and CMA-ES. We follow the officially released codebase and configurations of LatentDE, MLDE, TreeNeuralTS, TreeNeuralUCB, and EvoPlay for running. We take the results of PEX (one version with CNN model and another with MuFacNet model), AdaLead, CbAS, DbAS, and DyNA-PPO from Ren et al. (ren_proximal_2022) as we use the same benchmark setting, such as the starting sequence and fitness oracle. We run BO and CMA-ES with the implementations in FLEXS (sinai_adalead_2020). We average the results of AlphaDE, LatentDE, MLDE, TreeNeuralTS, TreeNeuralUCB, EvoPlay, BO, and CMA-ES over 5 independent trials. For all methods, we set the same oracle budget of 1000 queries. For AlphaDE, here we use the pretrained ESM2-series models (lin_evolutionary-scale_2023) for fine-tuning as it is one of the most powerful and widely used protein language models with various model sizes for study convenience. Specifically, in the standard AlphaDE, we use ESM2-35M for fine-tuning as its efficiency in model size and effectiveness in evolution plausibility. Details of the homologous sequence datasets for fine-tuning are in Appendix B, while hyperparameters of AlphaDE are given in Appendix C. For each task, the starting protein sequence for optimization is the sequence with the lowest protein fitness value, which is the same as the starting sequences in TAPE oracles. The benchmark results for these eight tasks are summarized in Table 1.
Table 1: Benchmarking in-silicon directed evolution methods. We present the maximum fitness in 1000 black-box oracle queries.
@@
LatentDE (tran_latentde_2025) 3.79 4.72 1.20 3.88 -3.34 -1.13 0.58 1.29
MLDE (tran_protein_2025) 2.04 -3.38 0.08 3.84 -1.38 -0.19 0.92 2.75
@@
The results in Table 1 show that AlphaDE significantly outperforms various baselines in most tasks. For example, in the task of AAV, the second best method MLDE obtains a fitness value of 4.72 while AlphaDE achieves 7.95 with an improvement of 68.43%. Similar situations also happen in the task of E4B and AMIE, which exhibit AlphaDE’s great evolution ability by combining effective searching (i.e., AlphaZero-like MCTS) with strong prior guidance (i.e., fine-tuned protein language model). At the same time, we also examine the diversity of evolved sequences by AlphaDE in Appendix I. AlphaDE with different ESM2 model sizes is tested in Appendix D. The ablation study of AlphaDE on the fine-tuning step and MCTS inference step is given in Appendix F. Benchmark results using another oracle based on ESM-1b further validate AlphaDE as shown in Appendix G. Moreover, the study of hyperparameter sensitivity of AlphaDE is available in Appendix H, while AlphaDE’s computation efficacy is provided in Appendix L.
At the same time, we also test AlphaDE with other protein language models such as ProtBert (elnaggar_prottrans_2022) (model size 420M) and ESM-1b (rives_biological_2021) (model size 650M), which are trained with the standard BERT (devlin_bert_2019) architecture using the UniRef100 (suzek_uniref_2007) database and with the optimized RoBERTa (liu_roberta_2019) architecture using the UniRef50 (suzek_uniref_2007) database, respectively. The results of AlphaDE with ProtBert and ESM-1b are shown in Table 2. We see that AlphaDE with ProtBert and AlphaDE with ESM-1b perform differently in some tasks, suggesting their evolution ability varies in different protein families. Generally, AlphaDE with the two different protein language models achieves superior performance, showing AlphaDE’s compatibility. Meanwhile, the fine-tuning step is the key to boosting performance, as directly using the pretrained versions of protein language models leads to a lot of fitness loss.